John bought a new fish tank with the measurement of 36 inch long 12 inches wide and 20 inches high he filled the tank 5 inches high with water he thenput in 8 fishes which made the water rise 3 and one half higher if he wants to leave 1and1half inches of space from the top of the tank how much more of the tank needs to be filled with water?
Answer by stanbon(48564)

The tank is 20" high.
It has 5" of water in it.
Fish add 1 1/2 " to the height of the water.
He leaves 1 1/2 " to spare.
To fill the rest of the height with water it needs to be
20" - (5 +3 1/2 +1 1/2)" = 20 - 10 = 10"
He need to raise the water 10 more inches.
Cheers,
Stan H>